### v3

# Leveraging LLMs for Top-Down Sector Allocation in Automated Trading

Abstract
This paper introduces a methodology leveraging Large Language Models (LLMs) for sector-level portfolio allocation through systematic analysis of macroeconomic conditions and market sentiment. Our framework emphasizes top-down sector allocation by processing multiple data streams simultaneously, including policy documents, economic indicators, and sentiment patterns. Empirical results demonstrate superior risk-adjusted returns compared to traditional cross-momentum strategies, achieving a Sharpe ratio of 2.51 and portfolio return of 8.79% versus -0.61 and -1.39% respectively. These results suggest that LLM-based systematic macro analysis presents a viable approach for enhancing automated portfolio allocation decisions at the sector level.
Preprint accepted to ICLR Workshop Advances in Financial AI: Opportunities, Innovations, and Responsible AI on March 5, 2025
1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed significant innovations in algorithmic trading and investment management, with numerous studies exploring their potential for stock selection and portfolio optimization (Ding et al., 2024). However, while existing research has predominantly focused on bottom-up security selection using company-specific metrics (see e.g., Yu et al. (2024); Zhang et al. (2024); Wang et al. (2024)), there remains a crucial gap in understanding how LLMs can enhance top-down investment strategies, particularly in the context of sector allocation within a portfolio.
Macroeconomic factors fundamentally shape investment decision-making, influencing stock valuations through both economic channels and market sentiment (Jareño and Negrut, 2016; Jabeen et al., 2022; Ma et al., 2024). A top-down investment approach, based on macroeconomic analysis, enables investors to anticipate broad market trends, optimize portfolio allocation across economic cycles, and identify systematic risks before they manifest in individual equity prices. These macroeconomic forces generate heterogeneous sectoral responses; for instance, inflationary pressures typically constrain manufacturing growth while benefiting services and agricultural sectors (Chaudhry et al., 2013). Moreover, hawkish Federal Reserve communications not only elevate borrowing costs and alter corporate decision-making but also intensify investor concerns about financially vulnerable firms, amplifying negative sentiment (Cieslak and McMahon, 2023).
This sentiment-driven reaction often precipitates market movements that exceed what fundamentals alone would otherwise suggest. Key macroeconomic variables—including monetary policy, inflation dynamics, and economic growth trajectories—propagate through the financial system, influencing both firms’ cost of capital and aggregate consumption patterns. Given that the task of constructing a portfolio based on sector allocation is one that requires piecing together information from different sources, this paper seeks to address the following Research Question – How can LLMs be leveraged to enhance top-down sector allocation strategies by integrating macroeconomic analysis with market sentiment for automated portfolio construction?
To respond to this question, we introduce a novel LLM-based methodology that emphasizes top-down sector allocation through integrated analysis of macroeconomic and market sentiment data. While existing approaches incorporate various data sources including sentiment and fundamentals, they typically focus on bottom-up analysis of individual securities. Our framework leverages LLMs to systematically process and synthesize multiple data streams simultaneously (including policy documents, economic indicators, and sentiment patterns), enabling dynamic adjustment of sector allocations based on market conditions. By automating the extraction and interpretation of these macro financial relationships, our framework enhances the responsiveness of sector allocation strategies through a predominantly top-down lens. This provides a more systematic approach to capturing sentiment-driven price movements, offering new insights into sector-level portfolio construction that complement traditional security-level analysis.
Results obtained through backtesting indicate that our sector allocation strategy significantly outperforms traditional cross-momentum approaches. Over the testing period, our methodology achieved a positive return of 8.79% and a Sharpe ratio of 2.51, compared to a loss of 1.39% and a Sharpe ratio of -0.61 for the cross-momentum strategy. These results suggest that our top-down LLM-based approach effectively captures sector-level opportunities while maintaining strong risk-adjusted performance, demonstrating the potential value of incorporating systematic macro and sentiment analysis in portfolio allocation decisions.
2 Related Works
In this section, we cover related literature on the following four topics: cross-sectional momentum investment, top-down investment, usage of LLMs as trading agents, and finance specfici Aspect based Sentiment Analysis.
2.1 Cross-sectional momentum investment
Traditional cross-sectional momentum strategies rank assets based on their relative performance, taking long positions in top performers and short positions in underperformers. The classical approach introduced by Jegadeesh and Titman (1993) ranks stocks based on past returns, while modern variations incorporate volatility normalization and machine learning models trained to minimize mean squared error before ranking assets to construct portfolios.
Learning to Rank (LTR) algorithms further improve upon traditional approaches by explicitly optimizing for ranking quality rather than prediction accuracy (Poh et al., 2020). These algorithms, which include pairwise methods like RankNet (Burges et al., 2005) and LambdaMART (Wu et al., 2010) and listwise methods like ListNet (Cao et al., 2007), learn the relative ordering between instruments directly. When applied to cross-sectional momentum strategies, LTR methods demonstrate significant improvements over traditional approaches, achieving approximately three times higher Sharpe ratios.
2.2 Top-down investment
The relationship between macroeconomic indicators and stock markets has been widely studied, demonstrating their significant impact on equity values.
Early research on Sri Lanka’s stock market (Gunasekarage et al., 2004) found that interest rates (measured by the treasury bill rate) and inflation (measured by the consumer price index) strongly influence stock prices, with the treasury bill rate exerting the most significant effect over a 17-year period (1985–2001).
More recent studies leverage macroeconomic regime modeling to enhance equity factor investing strategies (Nuriyev et al., 2024). Techniques such as Hidden Markov Models and meta-correlation-based clustering have outperformed traditional strategies by dynamically adjusting portfolio weights based on economic conditions, favoring quality factors during crises and momentum factors during growth phases. This further underscores the critical role of macroeconomic indicators in shaping stock market behavior.
2.3 LLM Trading Agents
Previous works (Ding et al., 2024) on LLM-based trading agents can be broadly categorized into (i) LLM as a Trader and (ii) LLM as an Alpha Miner. This literature review focuses on the former, examining works in News-Driven and Reflection-Driven LLMs as Traders.
2.3.1 LLM as a Trader - News Driven
News-driven LLMs integrate stock news and macroeconomic updates into their prompt context before instructing the LLMs to predict stock price movements.
Studies (Lopez-Lira and Tang, 2023; Wu, 2024) show that both proprietary and open-source LLMs, even without specialized financial training, can effectively analyze news sentiment for stock market prediction. Backtesting a simple long-short strategy based on sentiment scores from news headlines has demonstrated promising results.
A more advanced approach Fatouros et al. (2024) enhances the approach by summarizing and refining news data while reasoning about its relationship with stock price movements. A memory module stores these summaries, which are later retrieved as contextual ”recommendations” for trading decisions. By leveraging progressive daily news summaries, macroeconomic insights, and stock price momentum summaries, the MS-Top10-Cap-GPT strategy outperformed the S&P 100 index by approximately 30%.
2.3.2 LLM as a Trader - Reflection Driven
Apart from memories, reflections can be used in LLM decision making. In this context, Reflections (Park et al., 2023) are defined as high-level knowledge and insights progressively aggregated from raw memories and observations. The inclusion of memory and reflection in LLM-based algorithms offers significant benefits such as mitigating the risk of hallucinations (Ji et al., 2023) and obtaining high-level understanding of the environment (Park et al., 2023).
FinMem (Yu et al., 2024) introduces a trading agent with layered memorization and characteristics. The raw inputs, such as daily news and financial reports, are summarized into memories. Upon the arrival of new observations, the relevant memories are retrieved and integrated with these observations to produce reflections. Both memories and reflections are stored in a layered memory bucket. During the trading phase, these memories and reflections are retrieved and utilized by the decision-making module to generate the final trading decisions. The retrieval method considers the recency, relevancy, and importance of the information.
Similarly, FinAgent (Zhang et al., 2024) extends this approach with a multimodal module that processes numeric, text, and image data. It enhances decision making by incorporating technical indicators like MACD and RSI, along with analyst guidance, demonstrating superior backtesting performance over FinMem.
2.4 Finance specific Aspect based sentiment analysis
Traditional sentiment analysis in finance has historically focused on broad positive/negative classifications of text towards stocks, but markets and financial decisions require more nuanced understanding. Aspect based Sentiment Analysis (ABSA) has emerged as a crucial advancement by enabling the extraction and analysis of sentiment toward specific aspects or features within financial text Du et al. (2024); Malandri et al. (2018); Ma et al. (2023). Previous works of finance-specific ABSA (Ong et al., 2023) highlight the separation of sentiment across different financial aspects for explainable sentiment, allowing for assessment of sentiment towards specific risk factors or market aspects.
On the other hand, LLMs without finetuning were used for the task on finance specific ABSA. Through the Heterogeneous multi-agent Discussion (HAD) framework (Xing, 2024), the accuracy and F1 scores obtained on established datasets like FiQA (Maia et al., 2018), demonstrate the potential of LLMs to perform finance specific ABSA without specific financial training.
3 Data Source
This section provides an overview of the data sources utilized for this project. The backtesting period was limited to January 2019 through June 2019 due to constraints on computational resources and cost.
Our backtesting framework focused exclusively on S&P 500 constituents to maintain a well-defined set of liquid, large-cap stocks. To mitigate survivorship bias, we dynamically updated our universe to reflect real-time changes in the index composition. For instance, when First Republic Bank (FRC) replaced SCANA Corporation (SCG) following its acquisition by Dominion Energy on 2nd January 2019 (S&P Global, 2018), our set of stock used for backtesting was adjusted accordingly. Historical market data, including OHLC (open, high, low, close) prices and trading volumes, was sourced from the Alpha Vantage (2025) API.
The following Macro data were selected to be representative of the macroeconomic variables-interest rates, inflation, employment and economic growth-that represent the distinct economic categories as outlined by Nuriyev et al. (2024).
  • Inflation:
    • Consumer Price Index (CPI) (U.S. Bureau of Labor Statistics, 2025a)
    • Producer Price Index (PPI) (U.S. Bureau of Labor Statistics, 2025c)
    • Personal Consumption Expenditures (PCE) (U.S. Bureau of Economic Analysis, 2025)
  • Employment:
    • Non-Farm Payrolls (NFP) (U.S. Bureau of Labor Statistics, 2025b)
  • Economic Growth:
    • Purchasing Managers’ Index (PMI) (Investing.com, 2025)
  • Interest Rates:
    • Federal Open Market Committee (FOMC) minutes (Board of Governors of the Federal Reserve System, 2025)
Notably, the PCE data for 2018-Dec, 2019-Jan and 2019-Feb were released off schedule due to the partial federal government shutdown. As such, the latest available data from previous months was used instead.
By using keywords such as stock-specific company names and global market identifiers, news articles were scraped from the web via NewsAPI (2025) to be used as sentiment data. A total of 300000 articles were used to form the backtesting window of 6 months, 2019-Jan to 2019-Jun. Each news article consisted of the published date, title, description and content.
4 Design Process
Figure 1: Overview of agentic flow designed. Icons obtained from Flaticon.com
Top-down sector allocation investing requires a comprehensive understanding of both macroeconomic conditions and company-specific sentiments. As shown in our architecture diagram (refer to Fig 1), we address this through a dual-stream processing pipeline that analyzes news articles and macroeconomic data in parallel. The system integrates these inputs through a Ranking Agent which generates stock rankings that inform sector-level investment decisions. The architecture incorporates memory and reflection modules (represented as circles in the diagram) as discussed in Section 2.3.2 to help the system develop a high-level understanding of the macroeconomic environment.
As outlined in Section 2.4, recent research has demonstrated that LLMs can perform financial sentiment analysis without domain-specific training, guiding our design of the sentiment analysis pipeline. The Sentiment Agent, powered by the DeepSeek-7B-Chat model (DeepSeek-AI, 2025), processes news articles through two critical analyses: Named Entity Recognition (NER) and ABSA. NER is essential because our keyword search encompasses both stock-specific and broader market-relevant terms, enabling the identification of articles pertinent to target stocks that might be missed by simple keyword matching. The ABSA component analyzes article descriptions and content to identify multiple aspects of sentiment, providing the Ranking Agent with granular information about how macroeconomic conditions may amplify or diminish stock-specific sentiments through various channels. The resulting sentiment data (stock ticker and aspect sentiment pairs) is maintained in a dedicated Sentiment Memory Module, which serves as a persistent repository for historical sentiment information.
The DeepSeek-Chat-7B model, selected for three key advantages: its open-source nature, efficient inference speed, and cost-effective operation. The Sentiment Agent executes the analysis through a specialized prompt template (detailed in Appendix, Section 9.1).
Parallel to the sentiment pipeline, the macroeconomic data stream is processed by a Macro Agent. Since raw macroeconomic values lack contextual meaning for LLMs, this component calculates month-over-month percentage changes for key economic indicators including CPI, PPI, PCE, NFP, and PMI. These relative changes provide meaningful representations of macroeconomic trends that LLMs can effectively interpret and analyze. Additionally, the Macro Agent processes FOMC meeting minutes using DeepSeek-Chat-7B to generate concise summaries to be stored as memories in the Macro Memory Module for subsequent retrieval. For more information about the prompts used, please refer to the Appendix, Section 9.2.
The core of the agentic architecture is the Ranking agent, which retrieves memories from the News and Macro modules, as well as current portfolio positions from the Portfolio Management Agent, to generate a Stock Ranking reflection for a single day. To prevent look-ahead bias, the agent operates under strict temporal constraints, where it only processes news memories published on the previous trading day and macro memories based on their official release dates, using the latest available data points for each economic indicator. This temporal filtering mechanism ensures that all trading decisions are made using only information that would have been available to market participants at the time of trading, maintaining the practical applicability of our backtesting results.
The Ranking Agent implements a hierarchical top-down sector allocation strategy through three sequential analytical phases: (i) macroeconomic analysis, (ii) sector allocation, and (iii) sentiment integration. In the initial phase, the agent analyzes inflation trends, evaluates economic strength indicators, and assess monetary policy trajectories. This macroeconomic assessment helps identify helps which sectors perform the best in the given macroeconomic environment. The second phase focuses on portfolio optimization, where the agent determines optimal allocations across the 11 GICS sectors (S&P Global Market Intelligence, 2016) and evaluates necessary portfolio adjustments by comparing target allocations against current sector exposures. In the final phase, leverages sentiment-aspect pairs from the Sentiment Module to select stocks within each targeted sector. This refinement process evaluates how prevailing macroeconomic conditions may amplify positive sentiments or mitigate negative ones for specific business aspects, enabling the selection of stocks best positioned for both long and short positions within their respective sectors. For more information about the prompt used to generate the stock ranking reflections, please refer to Appendix, Section 9.3.
We employ the Llama-3.3-70B-Instruct model (Meta AI, 2024a) as the foundation for our Ranking Agent. This model selection was motivated by several key factors: its open-source nature, its extensive training on instruction-following tasks using publicly available datasets and over 25M synthetically generated examples. These characteristics make it particularly well-suited for interpreting and executing the complex, multi-phase analysis required for our sector allocation strategy.
The daily stock ranking reflection is processed by a Decision-making Agent, implemented using Llama-3.2-3B (Meta AI, 2024b), which converts qualitative rankings into structured JSON trading decisions. Given the standardized format of the reflections generated by the Ranking Agent, we opted for a lighter-weight instruction-following model to efficiently parse and transform these reflections into actionable trade signals. These decisions are then routed to the Portfolio Management Agent, which interfaces with real-time market data to optimize trade execution.
The Portfolio Management Agent implements a conservative capital management strategy designed to balance portfolio performance with risk mitigation. To maintain sufficient liquidity for opportunistic trading while preventing overexposure, the agent enforces a 90% maximum capital utilization threshold. During the execution of trades, new positions are initiated only when they do not create conflicts with existing holdings, whereas positions are fully liquidated when new rankings indicate a shift in stock sentiment outlook. Furthermore, to maintain alignment with our investment scope, positions in stocks removed from the S&P 500 index during the backtesting window are automatically liquidated.
5 Methodology
To evaluate the effectiveness of this architecture, we implement two distinct methodologies: our proposed top-down sector-based allocation strategy that leverages on LLM-driven macroeconomic analysis, and a baseline cross-momentum strategy that build a portfolio purely based on momentum factors.
Compared to the top-down sector allocation strategy, the ranking agent in the cross-momentum strategy would only carry out step (iii), integrating sentiment and macroeconomic information, to generate the reflection of stock rankings. Additionally, the position size of each trade is fixed at 1% of the portfolio. For more information about the prompt used, please refer to Appendix, Section 9.4.
An initial capital of USD100M was used to simulate real-world institutional trading conditions and provide sufficient scale to meaningfully assess the impact of transaction costs on strategy performance.
Trading costs were tracked through a comprehensive framework that accounts for both explicit and implicit costs. A commission rate of 10 basis points per trade was applied to reflect standard institutional brokerage fees, while market impact costs of 10 basis points were included to account for price slippage during execution.
6 Results
The portfolio percentage change and Sharpe ratio were selected as the primary metrics for evaluating the strategy, as they provide complementary perspectives on both performance and risk-adjusted returns. Portfolio percentage change offers a normalized view of performance, allowing for direct comparisons across different portfolio sizes and time periods by expressing gains and losses as a proportion of the initial investment. This facilitates a standardized assessment of the strategy’s performance. In contrast, the Sharpe ratio, which gauges excess returns relative to volatility, was chosen to evaluate the risk-adjusted efficiency of the strategy. Together, these metrics offer a comprehensive evaluation of the strategy’s profitability and its ability to generate sustainable returns.
The results presented in the Table 1 are calculated as follows: Portfolio percentage change is derived from the final portfolio value, which includes unrealized profits, minus trading costs. The Sharpe ratio is computed by dividing the daily excess returns by the standard deviation of the portfolio’s returns, providing a measure of risk-adjusted performance.
Table 1: Backtesting results from Jan-2019 to Jun-2019
Strategy PCT Change in Portfolio Sharpe Ratio
Cross-Momentum -1.39% -0.61
Sector-Allocation 8.79% 2.51
The results indicate that the Sector-Allocation strategy significantly outperformed the Cross-Momentum approach across both metrics. While the Cross-Momentum strategy resulted in a negative portfolio return of -1.39% and a Sharpe ratio of -0.61, suggesting poor risk-adjusted performance, the Sector-Allocation strategy delivered superior results with an 8.79% portfolio gain and a strong Sharpe ratio of 2.51. The positive Sharpe ratio for the Sector-Allocation strategy indicates that it not only generated better absolute returns but also achieved this performance with favorable risk-adjusted characteristics. This suggests that the sector-based approach was more effective at capturing market opportunities while managing volatility. The stark contrast in performance metrics between the two strategies highlights the benefits of incorporating top-down sector allocation in trading strategies.
7 Limitations and Future Directions
While our results demonstrate promising performance, several limitations and opportunities for future research warrant discussion. A primary limitation is the relatively short backtesting period from 2019 to 2024, constrained by computational resources and costs. Extending this period would provide more robust validation across different market cycles and economic conditions. Additionally, our current implementation could benefit from larger language models, which have demonstrated superior capabilities in complex Natural Language Processing tasks. Specifically, using models with more parameters could enhance the accuracy of ABSA, NER, and the interpretation of FOMC minutes by capturing more subtle linguistic nuances and relationships.
Several promising directions exist for future research. First, expanding the model’s data inputs to include company fundamental data, particularly quarterly earnings reports, could reveal whether incorporating bottom-up financial metrics alongside our top-down approach yields meaningful improvements in portfolio performance (Ong et al., 2025).
Second, the ranking mechanism could be enhanced through the implementation of LTR algorithms as mentioned in Section 2.1 or Reinforcement Learning techniques. These approaches have demonstrated success in improving cross-sectional strategies, suggesting potential benefits for sector allocation. By incorporating historical ranking performance as feedback, these methods could dynamically optimize allocation decisions and enhance the model’s adaptability to changing market conditions.
Additionally, integrating FINMEM’s (Yu et al., 2024) architecture could enhance the model’s ability to process multi-timeframe financial data while maintaining important information over extended periods, potentially capturing longer-term patterns and relationships that human traders might overlook.
Lastly, financial Explainable Artificial Intelligence techniques (Yeo et al., 2025) such as counterfactual explanations, can be used to evaluate performance in alternative scenarios. These methods provide insight into LLM decision making and support regulatory compliance.
8 Conclusion
In this paper, we introduced a novel methodology leveraging Large Language Models for top-down sector allocation through integrated analysis of macroeconomic and market sentiment data. Our approach demonstrates that LLMs can effectively process and synthesize multiple data streams simultaneously, enabling dynamic sector allocation adjustments based on market conditions. The empirical results validate the effectiveness of this approach, with our strategy achieving an 8.79% return and a Sharpe ratio of 2.51, significantly outperforming traditional cross-momentum approaches which recorded a -1.39% return and -0.61 Sharpe ratio during the same period.
These results suggest that systematic top-down analysis using LLMs offers a viable approach for enhancing portfolio allocation decisions. The framework’s ability to process unstructured data and identify macro financial relationships represents a step forward in automating complex market analysis tasks. As the field of quantitative investing continues to evolve, our methodology demonstrates the potential for LLMs to complement existing investment strategies by providing a more systematic approach to sector-level portfolio construction.
Acknowledgments
Gianmarco Mengaldo gratefully acknowledges the support of the MOE Tier 1 Startup project, under grant number #22-3565-A0001-1. The research is also supported by MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005) and RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore.